OpenStudy (anonymous):

The sum of the reciprocals of two consecutive even integers is 7/24. Write an equation that can be used to find the two integers. Find the two integers.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

If the integers are consecutive, we can call them x and x+1, their reciprocals means 1/x and 1/(x+1)... note the use of brackets, it's not 1/x+1 :P

OpenStudy (anonymous):

and then?

OpenStudy (agent0smith):

The sum means add them together \[\large \frac{ 1 }{ x}+\frac{ 1 }{ x+1 }=\frac{ 7 }{ 24 }\]

OpenStudy (dumbcow):

@agent0smith , its consecutive EVEN integers 1/x, 1/x+2

OpenStudy (anonymous):

So that's my answer but it doenst look like any of my answer choices

OpenStudy (agent0smith):

thanks @dumbcom \[\large \frac{ 1 }{ x}+\frac{ 1 }{ x+2 }=\frac{ 7 }{ 24 }\] you still have to solve that @SkiTTleoooo47 start by multiplying every term by x

OpenStudy (agent0smith):

@dumbcow dammit, another mistake lol

OpenStudy (anonymous):

okay no that looks like an answer choice but theres these numbers on the side 4 and 6 or 6 and 8?

OpenStudy (agent0smith):

you still have to solve. Multiply every term by the LCD 24x(x+2)\[\large 24(x+2)+24x=7x(x+2)\]now try expanding that and solving.

OpenStudy (anonymous):

How tf do i solve that lol

OpenStudy (agent0smith):

Expand. \[\large 24x+48+24x=7x^2+14x\]now combine like terms, and get everything on one side of the equals sign

OpenStudy (anonymous):

34x +48=7x^2

OpenStudy (agent0smith):

get everything on one side

OpenStudy (anonymous):

How would i put 7x^2 on the other side ?

OpenStudy (agent0smith):

Same way you got 14x on the other side.

OpenStudy (agent0smith):

Note that this does not give two consecutive even integers...it doesn't even have two integer solutions. http://www.wolframalpha.com/input/?i=%5Cfrac%7B+1+%7D%7B+x%7D%2B%5Cfrac%7B+1+%7D%7B+x%2B2+%7D%3D%5Cfrac%7B+7+%7D%7B+24+%7D one is a fraction. But i gotta go.